OpenStudy (ujjwal):

If the sum of the first n terms of an AP is \(nA+n^2B\) where A and B are constants, then the common difference of the AP is a)A+B b)A-B c)2A d)2B

OpenStudy (anonymous):

It is D..

OpenStudy (ujjwal):

How do you get it?

OpenStudy (anonymous):

See, if that is the sum of first n terms then: Put n= 1, you will get the first term: First term = A + B Now putting n= 2 you will get the sum of First two terms: Sum of First two terms : 2A + 4B Now subtract both of them you will get the second term: 2A + 4B - A - B = A + 3B This is the second term.. Now subtract second and first term you will get the common difference..

OpenStudy (foolaroundmath):

Sum of n-terms of an AP = \(\frac{n}{2}(2a + (n-1)d) \), a = first term, d = common difference \[S = n(a-\frac{d}{2}) + n^{2}\frac{d}{2}\] Comparing above equation with \(An + Bn^{2} \), we get \[A = a - \frac{d}{2}, B = \frac{d}{2}\] or \(d = 2B \)

OpenStudy (ujjwal):

Thanks @waterineyes and @FoolAroundMath

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

In any problem where sum is given to you in n form then find first and second term by this procedure and then find common difference...

OpenStudy (ujjwal):

yeah, i will be doing that now on wards! Thanks!!